Amalgams,blocks,weights, fusion systems and finite simple groups

Geoffrey R. Robinson, Department of Mathematical Sciences, University of Aberdeen, Aberdeen AB24 3UE, Scotland, g.r.robinson@abdn.ac.uk

ABSTRACT: We associate an iterated amalgam of finite groups to a certain class of fusion systems on finite p-groups (p a prime), in such a way that the pgroup of the fusion system is a maximal finite p-subgroup of the resulting group, unique up to conjugacy, and, furthermore, the conjugation action of the resulting (usually infinite) group on p-subgroups induces the original fusion system on the p-group. ^{1}. In view of earlier work of Puig and of Broto,Castellana, Grodal, Levi and Oliver, [5], the fusion systems we deal with include all saturated fusion systems.

The resulting amalgam has free normal subgroups of finite index, and we examine the images of the group by its maximal free normal subgroups of finite index; these images all contain (isomorphic copies of) the original p-group (and are generated by the (images of the) finite groups used in the amalgamation). If there is no non-trivial normal p-subgroup of the fusion system (equivalently, if the iterated amalgam constructed has no non-trivial normal p-subgroup), then the generalized Fitting subgroup of each of these homomorphic images is a direct product of non-Abelian simple groups, each of order divisible by p (and if there is no proper non-trivial strongly closed p-subgroup in the fusion system, then the generalized Fitting subgroup of any of these finite groups is characteristically simple).

We note that the (finite dimensional) representation theory of this amalgam is (almost by construction) p-locally determined. In the case of the fusion system associated to a p-block of a finite group, we suggest strong links between block-theoretic invariants of the above finite epimorphic images of the associated amalgam and of the original block. We believe that the results of this paper offer the prospect of at least the beginnings of a structural explanation for some of the current numerical conjectures in block theory (eg [2],[8]).

INTRODUCTION: Let p be a prime number and let P be a finite p-group. If P is contained in a (not necessarily finite) group G such that every finite psubgroup of G is G-conjugate to a subgroup of P, we will say that P is a Sylow p-subgroup of G (this notion is less general than that used by Aschbacher and Chermak in [4], but is sufficient for our present purpose). However, we warn the reader that even when P is a Sylow p-subgroup of G in the above sense, it is not necessarily the case that a proper subgroup of G need have a Sylow p-subgroup in the same sense (even if the subgroup is normal, so that arguments such as the Frattini argument are not necessarily available). It is perhaps worth remarking that when G has a Sylow p-subgroup, every subgroup, H of G has a unique largest normal finite p-subgroup, denoted (as usual) by O_{p}(H).

1Since this paper was written, we have been informed that I. Leary and R. Stancu are currently writing a paper with a different construction to realise a fusion system on a finite p-group via a group. Their work and ours are independent of each other

Finite group theorists used to working with local analysis will be familiar with the notion of determining a “fusion pattern” from some subgroup configuration assumed to be realised (typically) in a finite simple group. Similar considerations arise in block theory. This is formalized by the notion of a fusion system on a finite p-group P, introduced by L. Puig.

A fusion system on P is a category F_{P} (henceforth denoted F if there is no danger of ambiguity) whose objects are subgroups of P and whose morphisms are injective group homomorphisms from one subgroup of P into (but not necessarily onto) another satisfying a collection of axioms; these vary in the literature, but for our present purposes, key properties are the following:

i) All inner automorphisms of P are morphisms of F_{P} and Inn(P) is a Sylow p-subgroup of Mor_{F}(P, P). ii) If Q, R are subgroups of P, and ϕ ∈ Mor_{F}(Q, R), then the restricted group homomorphism Res_{U}^{Q}(ϕ) ∈ Mor_{F}(U, R) for each subgroup U of Q. If ϕ is surjective, then the group homomorphism ϕ^{−1} ∈ Mor_{F}(R, Q). If, furthermore, S, T are subgroups of P with S ≤ Qϕ and τ ∈ Mor_{F}(S, T), then the composite group homomorphism ϕ ∘ τ ∈ Mor_{F}(Q, T).

Whenever Q is a subgroup of P, and F_{Q} is a fusion system on Q with Mor_{F_{Q}}(S, T) ⊆ Mor_{F_{P}} (S, T) whenever S, T are subgroups of Q we say that the fusion system F_{Q} is contained in F_{P} .

We may also speak without ambiguity of the fusion sub-system on P generated by a collection of morphisms of F. It is the smallest fusion system on P which contains the given morphisms, all inner automorphisms on P, and is closed under composition, inversion (where applicable) and restriction of morphisms as above.

Notice that whenever Q is a subgroup of P, then Aut_{F}(Q) = Mor_{F}(Q, Q) is a subgroup of Aut(Q) which contains Inn(Q). The group Out_{F}(Q) is (as usual) defined to be Aut_{F}(Q)/Inn(Q). We say that Q is F-radical if O_{p}(Out_{F}(Q)) = 1.

If Q and R are subgroups of P such that Mor_{F}(Q, R) and Mor_{F}(R, Q) are both non-empty, then we say that Q and R are F-isomorphic, or sometimes F-conjugate.

We say that Q ≤ P is F-centric if C_{P} (Qϕ) = Z(Qϕ) whenever ϕ is a morphism of F whose domain contains Q. Notice that overgroups (in P) of F-centric subgroups are also F-centric. A subgroup Q of P is said to be fully F-normalized if |N_{P} (Q)| ≥ |N_{P} (Qϕ)| for all morphisms ϕ of F whose domain contains Q. A subgroup Q of P is said to be fully F-centralized if |C_{P} (Q)| ≥ |C_{P} (Qϕ)| for all morphisms ϕ of F whose domain contains Q. A subgroup V of P is said to be strongly closed in P with resepect to F if Xϕ ≤ V for each subgroup X of V and each ϕ ∈ Mor_{F}(X, P).

Puig also introduced the notion of a saturated fusion system, in an attempt to encapsulate the essential features arising in fusion systems from finite group theory and block theory. Once more, there are several equivalent formulations of this notion in the literature, and we do not need to give an explicit definition here, since the substitute notion of an Alperin fusion system (which includes saturated fusion systems) suffices for our purposes.

Loosely speaking, saturated fusion systems are modelled on the fact that for a given finite group, the p-Sylow theorems hold in the normalizers and centralizers of p-subgroups.

We will be primarily interested here in fusion systems on P in which the morphisms are precisely those which are induced by conjugations in a (not necessarily finite) group G which have P as a Sylow p-subgroup in the earlier sense. We will refer to this as the G-fusion system on P and say that the given fusion system on P is realised by G. In the case of finite groups, the full strength of Sylow’s theorem imposes extra conditions on fusion systems, which ultimately lead to Alperin’s fusion theorem [1] and its stronger variants due to Goldschmidt [9] and Puig [13]. In fact, the conditions on a saturated fusion system are such that (interpreted suitably) Alperin’s fusion theorem (and its refinements) hold for saturated fusion systems, as noted in Broto,Grodal,Castellana, Levi and Oliver (henceforth sometimes referred to as BCGLO for brevity) [5].

We will say that the fusion system F_{P} = F is an Alperin fusion system if there are subgroups E_{1}, E_{2}, … E_{r} of P and (finite) groups L_{1}, …, L_{r} such that:

i) E_{i} = F^{∗}(L_{i}) for each i (equivalently, E_{i} = O_{p}(L_{i}) (the largest normal psubgroup of L_{i}) and C_{L_{i}}(E_{i}) = Z(E_{i})). ii) L_{i}/E_{i} ≅ Out_{F}(E_{i}) for each i. iii) P_{i} = N_{P} (E_{i}) is a Sylow p-subgroup of L_{i} for each i and P_{1} = P. iv) For each i, the L_{i}-fusion system on P_{i} is contained in the fusion system F_{P} .

v) The fusion system F_{P} is generated by all the fusion systems of the L_{i} on the respective P_{i}.

Alperin’s fusion theorem and its various later refinements show that when P is a Sylow p-subgroup of a finite group G, the G-fusion system on P is an Alperin fusion system in the above sense. However, even in this case, the L_{i}’s need not be subgroups of G. Results of Alperin-Broué [3] and Külshammer-Puig [11] yield that if P is a defect group for a p-block B, then the fusion system on P induced by the conjugacy of B-subpairs is an Alperin fusion system in the above sense.

It is a consequence of 4.3 of Broto,Castellana, Grodal, Levi and Oliver, [5], that all saturated fusion systems are Alperin fusion systems in the sense above, the key point to note ( for our purposes) being the existence of an extension of Out_{F}(E_{i}) by E_{i} with the “correct” Sylow p-subgroup and fusion system, this extension being unique subject to these properties.

Broto, Levi and Oliver (sometimes with coauthors Grodal and Castellana) have studied abstract fusion systems on finite p-groups from a homotopy-theoretic viewpoint, introducing the concept of p-local finite groups (eg [6],[7],[12]).

There are finite p-groups P having an Alperin fusion system F_{P} which may not be realised by any finite group G. Example 3.4 of Levi and Oliver, [13], shows that the Solomon fusion system is one such system, since this fusion system is known to be saturated. In the same work, Aschbacher and Chermak constructed an infinite group G which realises this fusion system.

Aschbacher and Chermak used amalgams in their constructions. Our constructions below also use amalgams in an essential way. The key properties of the amalgam construction and its relationship to fusion which we make use of here are in essence (and inevitably) the same properties as those used in [4], though we regard our approach as more direct and our main construction is an iterative one involving forming several (though a finite number) of amalgams. The iterated amalgam we exhibit could also be identified with a direct limit of a finite collection of finite groups. However, constructing it iteratively makes its fusion system more transparent and also illustrates that the direct limit does not collapse.

Another point of departure from the work of Aschbacher and Chermak (and most previous authors) is that we amalgamate over p-subgroups only. This is (in the first instance at least) a naive viewpoint, but discarding some of the available identifications ultimately proves to be an advantage. Furthermore, there is no reason to insist that the groups we amalgamate have a common Sylow p-subgroup.

The idea of using amalgams and direct limits to “manufacture” groups with desired conjugacy properties has a long history, (including, for example, the HNN-construction of G. Higman, B.H. Neumann and H. Neumann, and the construction of infinite groups in which all elements of the same order are conjugate). In any event, such results illustrate that the category of finitely generated groups is large enough to exhibit groups with (what might be considered to be) pathological conjugacy properties; this can be turned to advantage when attempting to exhibit groups which realise a prescribed fusion system on a finite p-group.

Our main aim here is to prove:

THEOREM 2: Let F_{P} = F be an Alperin fusion system on a finite p-group P and associated groups L_{1}, L_{2}, …, L_{r} as in the definition. Then there is a finitely generated group G which has P as a Sylow p-subgroup such the fusion system F is realised by G. Furthermore, G may be chosen so that:

i) G has a free normal subgroup K of finite index and if r > 1, then any finite normal subgroup of G is a p-group. ii) Whenever Q is an F-centric subgroup of P, the group C_{G}(Q)/Z(Q) is torsionfree, C_{G}(Q) has a free normal subgroup of finite index, and all subgroups of C_{G}(Q) which intersect Z(Q) trivially are free. iii) If there is no non-trivial subgroup of P which is normal in each of the groups L_{i} (equivalently, if O_{p}(G) = 1), and K of part i) is chosen to be a maximal free normal subgroup of finite index, then F^{∗}(G/K) is a direct product of non-Abelian simple groups, and each minimal normal subgroup of G/K has nontrivial intersection with P K/K. If, furthermore, no proper non-trivial subgroup of P is strongly closed in P with respect to F, then G/K has a unique minimal normal subgroup.

REMARK: Some comments are in order. In general, it need not be the case in part iii) that K (or even [G : K]) is unique. Also, it is not always the case that P K/K is a Sylow p-subgroup of G/K. We remark also that while the work of B(CG)LO is directed towards associating a p-completed classifying space to a saturated fusion system, an issue, which is not, as far as we can see, resolved by our present work, from a group-theoretic and block-theoretic point of view, we believe that the group G and its finite homomorphic images as in part iii) will provide adequate substitutes for many purposes. For example (except in the degenerate case when some non-trivial subgroup of P is normal in each L_{i}), Theorem 2 does naturally associate finite simple groups to any Alperin fusion system. In the case of groups or blocks, the degenerate case corresponds to the case when fusion is controlled within the normalizer of a single subgroup or subpair respectively.

In the definition of an Alperin fusion system, there may be much redundancy. In the familiar types of fusion system (those arising from conjugacy in finite groups, from block theory, or a general saturated fusion system) this redundancy can be eliminated in the following sense: in these cases, there is a natural minimal choice of the L_{i}’s and E_{i}’s, an observation of Puig [13], building on earlier work of Goldschmidt [9], known as the “essential” subgroups. These have the additional property that we may label so that L_{1}/E_{1} either has p′- order, or has a strongly p-embedded proper subgroup and for i > 1, L_{i}/E_{i} has a strongly p-embedded proper subgroup (and we only need one E_{i} from each F-conjugacy class of such subgroups).

If we take the Alperin fusion system of a finite simple group of Lie type and rank greater than 1 in characteristic p, which is generated by the parabolic subgroups containing a given Borel, then the essential collection of L_{i}’s consists precisely of the minimal parabolics containing that Borel. This is a situation where even the Sylow normalizer is redundant.

In general, we could insist that the fusion system F_{P} is generated by the fusion systems of the L_{i}’s on the P_{i}’s but not by the fusion systems of any proper subcollection of the L_{i}’s on the P_{i}’s ( we can simply discard L_{i}’s until this situation is achieved). We say that the Alperin fusion system is irredundant in that case.

In general, we consider that (at least in the saturated case), the construction of the group G and its finite epimorphic images as in part iii) of Theorem 2 offer at least the suggestion of a beginning for a general “Lie theory” for any fusion system, a theory which Alperin suggests in [2] should exist for blocks; in particular, there should be such a theory for any prime, for any finite group (and for any block). The essential L_{i}’s ( for i > 1) play the role of minimal parabolics ( though, in general, they need not all contain P). We intend to expand on these comments in a future paper.

When k is the algebraic closure of GF(q) (for some prime q, possibly q = p) G is as in Theorem 2, and V is a finite-dimensional kG-module, then (since G is finitely generated and k is a locally finite field), we see that G/C_{G}(V) is a finite group (without non-trivial normal p-subgroup if V is irreducible and q = p). The finite-dimensional k-representation theory of G is determined (because of the amalgam construction) by the representation theory of the L_{i}’s in the sense that, given a “candidate” kG-module U, on which each L_{i} acts, we only need to check that Res_{P_{i}}^{P}(U) yields the same action of P_{i} as Res_{P_{i}}^{L_{i}}(U) for each i to know that U has a genuine kG-module structure. We will return to this issue in more detail in a later section.

We remark also that when q = p and we have a kP -module V with a group homomorphism ϕ_{i} : L_{i} → Aut(End_{k}(V)) for each i such that each ϕ_{i} restricts to P_{i} in a fashion compatible with the action of P on V, then we obtain an irreducible projective representation of G (in Schur’s sense).

In a more speculative mode, and continuing the spirit of some of our earlier work in block theory, e.g. the paper [10] by Kessar, Linckelmann and the present author, and [15], [16], which we consider indicates that natural numerical block theoretic invariants should be determined by the fusion of the block, and certain associated 2-cocycles, we suggest the following:

CONJECTURE: Let B be a p-block with defect group P of a finite group H. Let F_{B} = F be the fusion system on the B-subpairs ( with the H-conjugation providing morphisms). Then one of the following occurs:

i) There is a B-subpair (Q, b_{Q}) ⊲ N_{H}(P, b_{P}) with Q ≠ 1 such that, setting B^{∗} to be the unique block of N_{H}(Q, b_{Q}) lying over b_{Q}, there is a defect preserving bijection between complex irreducible characters of B and complex irreducible characters of B^{∗}, such that, furthemore, we have ℓ(B) = ℓ(B^{∗}). ii) Letting G be constructed from F as in part iii) of Theorem 2, there is at least one choice of free normal subgroup K of G such that, setting H^{∗} = G/K, there is a block B^{∗} with defect group (naturally identified with) P^{∗} of a central extension of H^{∗} by a cyclic p′-group such that there is a defect preserving bijection between complex irreducible characters of B and complex irreducible characters of B^{∗}, such that, furthemore, we have ℓ(B) = ℓ(B^{∗}).

A crucial step in the proof of Theorem 2 is the rather general Theorem 1 below: this is a procedure whereby a fusion system on a finite p-group P may be amalgamated with a fusion system on a subgroup Q of P, when these are realised by groups G and H respectively. The resulting “amalgamated” fusion system on P is realised by the amalgam G ∗_{Q} H.

Our result here is:

THEOREM 1: Let G be a group which has a Sylow p-subgroup P and H be a group which has a Sylow p-subgroup Q for some subgroup Q of P. Then the group K = G ∗_{Q} H has a Sylow p-subgroup P, and the K-fusion system on P is precisely that generated by the G-fusion system on P and the H-fusion system on Q.

SECTION 1: GENERAL PROPERTIES OF AMALGAMS

Let A, B be groups with a common subgroup C. Let G = A ∗_{C} B, where the amalgam is formed using the two inclusions. Then A and B may be identified with subgroups of G so that A ∩ B = C. For the basic facts about the structure of amalgams which we may make use of without further reference, we refer the reader to Serre [17].

The next Lemma allows us to pinpoint precisely how G-conjugacy of subgroups of A and B is effected.

LEMMA 1: Let X be a subgroup of A and g ∈ G\A such that X^{g} ⊆ A or X^{g} ⊆ B. Write (as we may) g = a_{0}b_{1}a_{1} … b_{s}a_{s}b_{∞}, where a_{i} ∈ A\C for 1 ≤ i ≤ s, b_{i} ∈ B\C for 1 ≤ i ≤ s, a_{0} ∈ A, b_{∞} ∈ B. Set X_{0} = X^{a_{0}}, Y_{1} = X_{0}^{b_{1}}, X_{1} = Y_{1}^{a_{1}}, Y_{2} = X_{1}^{b_{2}}, etc. .Then ⟨X_{0}, X_{i}, Y_{i} : 1 ≤ i ≤ s⟩ ⊆ C.

PROOF: The form of reduced words in G tells us that a product w of elements of G whose successive terms are in alternate members of {A, B}, and which all lie outside C, itself lies outside C. Hence if such a product has length greater than 1, it even lies outside either of A or B (for example, if the product w equates to the element a of A\C, we may suppose that the rightmost term of the given product expression for w lies in B\C, in which case the extended product expression wa^{−1} can’t represent the identity (certainly an element of C)).

We may multiply g on the right by an element of C if necessary, so we may suppose that either b_{∞} ∈ (B\C) or b_{∞} = 1. In either case, if there is some element u ∈ X^{a_{0}}\C, the fact that

b_{∞}^{−1} … a_{1}^{−1}b_{1}^{−1}ub_{1}a_{1} … b_{s}a_{s}b_{∞}

lies in A or B contradicts the result just established. The same argument produces a contradiction if a_{1} is present and X_{0}^{b_{1}} is not contained in C. We now proceed inductively and arrive at the stated conclusion.

REMARKS: i) Lemma 1 allow the use of amalgams to construct groups G in which Sylow’s theorems fail, even when all finite p-subgroups of G have bounded order. For if G = A ∗_{C} B is an amalgam as above, we conclude that a subgroup X of A which is not A-conjugate to a subgroup of C can’t be G-conjugate to a subgroup of B. It is also known (see Theorem 4.3.8 of [17]) that every finite subgroup of G is conjugate to a subgroup of A or of B. So if ( for any prime p), we take A and B to be cyclic of order p and C to be trivial (so that G is just the free product of two groups of order p) we see that A and B are maximal finite p-subgroups of G which are not conjugate in G.

ii) We also note that the Lemma places restriction on normalizers and centralizers of subgroups of A (and likewise, by symmetry, B). For when X ≤ A with no A-conjugate of X contained in C, we conclude that N_{G}(X) ≤ A.

Our next Lemma will be used later to exhibit free normal subgroups of finite index of amalgams we will construct. It is an algebraic substitute for some geometrical and topological constructions.

LEMMA 2: Let A, B be groups with a common finite subgroup C. Let H be a subgroup of A containing C. Suppose that there are homomorphisms ϕ : A → S_{n} and ψ : B → S_{m} with respective kernels K, L such that that H ∩K = C ∩L = 1. Then (identifying H with the corresponding subgroup subgroup of A ∗_{C} B) there is a homomorphism θ : A ∗_{C} B → S_{lcm([A:K],[B:L])} such that H ∩ kerθ = 1.

PROOF: We may embed A/K in S_{[A:K]} by using its regular representation. Similarly, we may embed B/L in S_{[B:L]}. In these new permutation representations, H and C respectively are acting freely ( or semi-regularly in group theoretic terminology). Setting ℓ = lcm([A : K], [B : L]), and taking a direct sum of ℓ/[A : K] copies of the first permutation representation and ℓ/[B : L] copies of the second gives us homomorphisms α : A → S_{ℓ} and β : B → S_{ℓ} which may be taken to agree on C. By the universal property of the amalgam, there is a common extension of α and β to a homomorphism θ : A ∗_{C} B → S_{ℓ} and clearly H ∩ kerθ = 1 as H acts freely.

SECTION 2: PROOFS OF THEOREMS

PROOF OF THEOREM 1: We set K = G ∗_{Q} H (the sum of G and H with Q amalgamated, in the terminology of [17]). By (the corollary to) Theorem 4.3.8 of [17], ( and using Sylow hypotheses on P, Q in G, H respectively), we see that every finite p-subgroup of K is conjugate to a subgroup of P; in other words, P is a Sylow p-subgroup of K in the earlier sense.

Identifying G and H with subgroups of K, we note immediately that the K-fusion system on P includes that generated by the G-fusion system on P and the H-fusion system on Q. We need to establish that it contains no more than this.

By Lemma 1, if X is a subgroup of G, h_{1}, …, h_{r+1} are elements of H\Q and g_{1}, …, g_{r} are elements of G\Q such that X^{h_{1}g_{1}...h_{r}g_{r}h_{r+1}} is contained in G, then all of the intermediate subgroups X, X^{h_{1}}, X^{h_{1}g_{1}}, …, X^{h_{1}g_{1}...h_{r}g_{r}h_{r+1}} are contained in Q (we allow the possibility that r = 0, in which case no element of G appears in the product, and a single element of H).

Now let X be a non-trivial subgroup of P, and suppose that ⟨X, X^{u}⟩ ⊆ P for some u ∈ K. We need to show that conjugation by u is in the fusion system generated by the fusion system of G on P and the fusion system of H on Q. If u ∈ G, there is nothing to do, so suppose that this is not the case. Then there are (possibly identity) elements g_{−∞}, g_{∞} ∈ G and elements h_{1}, g_{1}, …, h_{r}, g_{r}, h_{r+1} with each g_{i} ∈ G\Q and each h_{i} ∈ H\Q (again, r = 0 is allowed, in which case it is understood that u = g_{−∞}h_{1}g_{∞}) such that u = g_{−∞}h_{1}g_{1} … h_{r}g_{r}h_{r+1}g_{∞}.

Let X^{∗} = X^{g_{−∞}}. Then ⟨X^{∗}, (X^{∗})^{h_{1}g_{1}...g_{r}h_{r+1}}⟩ is contained in G, so that (by Lemma 1) all the intermediate groups X^{∗}, (X^{∗})^{h_{1}}, (X^{∗})^{h_{1}g_{1}}, …, (X^{∗})^{h_{1}g_{1}...h_{r}g_{r}h_{r+1}} are contained in Q.

Now (if present), the conjugation from X to X^{∗} is effected within the fusion system of G on P (since it is effected via conjugation by g_{−∞} ∈ G). The conjugation from X^{∗} to (X^{∗})^{h_{1}} is effected within the fusion system of H on Q, and if r > 0, for 1 ≤ i ≤ r, the conjugation from (X^{∗})^{h_{1}...h_{i}} to (X^{∗})^{h_{1}...h_{i}g_{i}} is effected within the fusion system of G on P ( since both the originating subgroup and the terminating subgroup are contained in Q, hence in P, and the conjugating element is in G) while the conjugation from (X^{∗})^{h_{1}...h_{i}g_{i}} to (X^{∗})^{h_{1}...h_{i}g_{i}h_{i+1}} is effected within the fusion system of H on Q ( since both the originating subgroup and the terminating subgroup are contained in Q, and the conjugating element is in H).

Finally, (if present), the conjugation from (X^{∗})^{h_{1}...h_{r}g_{r}h_{r+1}} to (X^{∗})^{h_{1}...h_{r}g_{r}h_{r+1}g_{∞}} is effected within the fusion system of G on P ( since the originating subgroup is contained within Q, hence in P, while the terminating subgroup (which is X^{u}) is contained in P, and the conjugating element is in G). This completes the proof of Theorem 1.

PROOF OF THEOREM 2: Given that F = F_{P} is an Alperin fusion system via the associated groups L_{1}, … L_{r}, with respective normal p-subgroups E_{1}, E_{2}, … E_{r}, and respective Sylow p-subgroups P_{i} (each contained in P = P_{1}), we inductively define groups K_{i} which (via Theorem 1) each have P as a Sylow p-subgroup as follows: K_{1} = L_{1} and for 1 ≤ i ≤ r, we set K_{i+1} = K_{i} ∗_{P_{i+1}} L_{i+1}. We set G = K_{r}.

Repeated applications of Theorem 1 yield that the G-fusion system on P is precisely F_{P} . We claim that whenever Q ≤ P is F-centric, C_{G}(Q)/Z(Q) is a torsion free group. First of all, if y ∈ C_{G}(Q) has finite order, then Q⟨y⟩ is contained up to conjugacy in some L_{i}. Hence we may suppose that Q ≤ P_{i} for this i and that y ∈ C_{L_{i}}(Q). Notice that C_{P_{i}}(Q) = Z(Q) since Q is F-centric. Replacing Q by an L_{i}-conjugate if necessary so that P_{i} contains a Sylow psubgroup of QC_{L_{i}}(Q), and using the fact that F^{∗}(L_{i}) = E_{i}, a now standard argument yields that C_{L_{i}}(Q) is a p-group, so, in fact C_{L_{i}}(Q) = Z(Q).Hence we have y ∈ Z(Q). Thus C_{G}(Q)/Z(Q) is torsion free. Since G is an iterated amalgam of finite groups, all its torsion free subgroups are free. Any finite normal subgroup N of G is conjugate to ( hence contained in) some L_{i}. If r > 1, then by the remark following Lemma 1, N must be a p-group, since G = N_{G}(N).

For part i) note that, by Lemma 2, there is a normal subgroup K of G with K ∩P = 1 and [G : K] finite. Since all elements of order p in G are conjugate to elements of P and all elements of finite order in G are contained in a conjugate of one or more of the L_{i}’s, we see that K ∩ L_{i} is a (finite) normal p′-subgroup of L_{i} for each i, so is trivial since F^{∗}(L_{i}) is a p-group for each i. Hence K is free, by 4.3.8 of Serre [17].

If we assume that K is maximal subject to being normal and free of finite index, then whenever K < N ⊲ G, we must have P ∩ N ≠ 1, otherwise the argument just used for K yields that N is torsion free, and then free. We may suppose that N/K is a minimal normal subgroup of the finite group G/K. Now

1 ≠ (P ∩ N) ≅ (P ∩ N)/(P ∩ K) ≅ K(P ∩ N)/K ≤ N/K,

so that N/K is not a p′-group. Suppose that N/K is an elementary Abelian p-group, and set U = P ∩ N ⊲ P, which is also elementary Abelian (and is non-trivial).

Choose any one of the subgroups L_{i} given by the Alperin fusion system. For ease of notation, set L = L_{i} and E = O_{p}(L)(= E_{i}).

Now

N ∩ L ≅ (L ∩ N)/(L ∩ K) ≅ K(L ∩ N)/K ≤ N/K,

so that L ∩ N is an elemenetary Abelian p-group. Since N ⊲ G, we see that L ∩ N ≤ E = O_{p}(L). Since G realises the fusion system F, we have

N_{G}(E)/EC_{G}(E) ≅ Out_{F}(E) ≅ L/E.

Since

(L ∩ EC_{G}(E)) = EC_{L}(E) = E,

we see that N_{G}(E) = LC_{G}(E).

We claim that U ⊲ L. It suffices to prove that U ≤ E, for then

U = P ∩ N ∩ L = P ∩ E ∩ N ∩ L = E ∩ N ∩ L ⊲ L.

Since EU is a finite p-group, it suffices to prove that N_{U}(E) ≤ E, for then N_{EU}(E) = EN_{U}(E) ≤ E, so E = EU.

Since U is an Abelian normal subgroup of P, we know that [E, N_{U}(E)] ≤ (U ∩ E) ≤ (L ∩ N) ⊲ L. Now

[L ∩ N, N_{U}(E)] = [L ∩ N ∩ P, N_{U}(E)] ≤ [U, N_{U}(E)] = 1.

Hence N_{U}(E) centralizes the normal series of p-subgroups 1 ≤ (L ∩ N) ≤ E of N_{G}(E) = LC_{G}(E). Thus (EN_{U}(E)C_{G}(E))/EC_{G}(E) is contained in O_{p}(N_{G}(E)/EC_{G}(E)) = 1.

Thus

EN_{U}(E) ≤ (EC_{G}(E) ∩ EU) = EC_{UE}(E) ≤ EC_{P} (E) = E,

as required.

Thus we have shown that every minimal normal subgroup of G/K is a direct product of non-Abelian simple groups, each of order divisible by p.

Notice that whenever M is a normal subgroup of G strictly containing K, then V = M ∩ P ≠ 1 by the choice of K, and V is strongly closed in P with respect to G ( for the benefit of the reader, we recall that this means that V^{g} ∩ P ≤ V for all g ∈ G). Hence if no proper non-trivial subgroup of P is strongly closed in P with respect to G, we conclude that G/K has a unique minimal normal subgroup, which contains P K/K.

SECTION 3: MODULES FOR FUSION SYSTEM AMALGAMS

Let F = F_{P} be an Alperin fusion system with associated groups L_{1}, L_{2}, …, L_{r}. For each i let P_{i} be a Sylow p-subgroup of L_{i} with P_{i} ≤ P = P_{1} for each i. Set E_{i} = F^{∗}(L_{i}) = O_{p}(L_{i}) for each i.

Let G be the iterated amalgam

(… (L_{1} ∗_{P_{2}} L_{2}) ∗_{P_{3}} L_{3}) ∗ …) ∗_{P_{r−1}} L_{r−1}) ∗_{P_{r}} L_{r}.

Then we see inductively that given a group M and group homomorphisms ϕ_{i} : L_{i} → M for 1 ≤ i ≤ r with

Res_{P_{i}}^{L_{i}}(ϕ_{i}) = Res_{P_{i}}^{P_{1}}(ϕ_{1})

for each i, there is a unique group homomorphism ϕ : G → M which extends each ϕ_{i} ( this justifies our earlier statements about G being identified with a direct limit).

We wish to begin the study the finite-dimensional kG-modules. Notice that if H is a finite group generated by subgroups M_{i} : 1 ≤ i ≤ r such that for each i there is a group epimorphism α_{i} : L_{i} → M_{i} with Res_{P_{i}}^{P_{1}}(α_{1}) = Res_{P_{i}}^{L_{i}}(α_{i}), then H is an epimorphic image of G.

We allow the possibility that U = O_{p}(G) ≠ 1. We remark that U ≤ E_{i} for each i, since

N_{U}(E_{i})E_{i}C_{G}(E_{i})/E_{i}C_{G}(E_{i}) ≤ O_{p}(N_{G}(E_{i})/E_{i}C_{G}(E_{i}) ≅ O_{p}(Out_{F}(E_{i})) = 1.

THEOREM 3: For each i, let H_{i} be a p′-subgroup of L_{i}, and let t = lcm{[L_{i} : H_{i}] : 1 ≤ i ≤ r}. Then there is a group homomorphism ϕ : G → S_{t} whose kernel is a free group.

PROOF: By a now familiar argument, it is enough to construct the homomorphism ϕ so that kerϕ has trivial intersection with P. Let Ω = {1, 2, …, t} and let each L_{i} act as it would on the direct sum of t/[L_{i} : H_{i}] copies of the permutation module of L_{i} on the cosets of H_{i}. Since P = P_{1} acts semi-regularly, and each P_{i} does, we may label the points so that each P_{i} acts in the manner determined by regarding it as a subgroup of P_{1}. By the remarks preceding the Theorem, Ω now has the structure of a G-set. Since the action of P is free, the kernel of the action is a free normal subgroup of finite index.

REMARKS: We will see presently that the transparent situation above is typical of the case when we have a collection of projective modules (in characteristic p) for the respective L_{i}. However, in general, there is no need to insist on free action of the P_{i}, we only need a faithful action of P which restricts to the action of the P_{i} on the given L_{i}-module ( or L_{i}-set). The following example may be illuminating, and illustrates that P K/K need not be a Sylow p-subgroup of G/K in Theorem 2.

EXAMPLE: Let P be a dihedral group of order 16. If we consider the 2-fusion in a group H ≠ O^{2}(H) with P as a Sylow 2-subgroup, then the L_{i}’s we obtain are L_{1} = P and L_{2} = S_{4}. Hence G may be taken to be the amalgam S_{4} ∗_{D_{8}} P in this case.

It is shown in [17] That for every prime p ≡ ±1 (mod 16), the group PSL(2, p) occurs as an epimorphic image of G (and that the kernel K = K(p) is a torsion free normal subgroup, hence free as in earlier arguments). Also, the explicit realization of G as a subgroup of SO(3, ℝ) given there easily yields that N_{G}(P) = P. For appropriate choices of p, the Sylow 2-subgroup of PSL(2, p) may be chosen to be dihedral of any specified order greater than 16.

In that case, K is a maximal normal subgroup of G, and setting \overline{G} = G/K, we see that [N\overline{G}(\overline{P}) : \overline{P} C\overline{G}(\overline{P})] is even, so that certainly \overline{N_{G}(P)} ≠ N\overline{G}(\overline{P}). We must conclude that KP has more than one conjugacy class of maximal 2-subgroups, and that the Frattini argument is not available in N_{G}(KP).

We now let k denote the algebraic closure of GF(p). We next mention a remarkable representation-theoretic property of G. When V is a kG-module, we let C_{G}(V) denote the kernel of the action of G on V. As remarked earlier, [G : C_{G}(V)] is finite if V is finite-dimensional, since k is a locally finite field and G is finitely generated.

THEOREM 4: i) For 1 ≤ i ≤ r let X_{i} be a finite-dimensional projective kL_{i}-module, and suppose that all X_{i} have equal dimension. Then there is a kG-module X such that Res_{L_{i}}^{G}(X) ≅ X_{i} for each i. Furthermore, C_{G}(X) is a free normal subgroup of G of finite index.

ii) For each i, let V_{i} be a simple kL_{i}-module. Then there is a finite dimensional simple kG-module V such that for each i, soc(Res_{L_{i}}^{G}(V)) is isomorphic to a direct sum of copies of V_{i}.

PROOF: i) Since X_{i} is a free kP_{i}-module for each i, we may suppose that P_{i} has the same action on X_{i} as it does on Res_{P_{i}}^{P_{1}}(X_{1}) for each i. In that case, there is a unique way to extend the action of the L_{i} on the underlying k-vector space to an action of G on that space. We let X denote the kG-module so obtained. Then X is a free kP -module, so that C_{G}(X)∩P = 1 and C_{G}(X) is a free normal subgroup of G of finite index.

ii) Let Y_{i} denote the projective cover of V_{i} as kL_{i}-module, so that we also have soc(Y_{i}) ≅ V_{i}. Let ℓ = lcm{dim_{k}(Y_{i})}. For each i, let X_{i} be a direct sum of ℓ/dim_{k}(Y_{i}) copies of Y_{i}, and let X be a kG-module constructed from the X_{i} as in part i). Then for each i, soc(Res_{L_{i}}^{G}(X)) is isomorphic to a direct sum of copies of V_{i}. Let V be a simple kG-module of X. Then soc(Res_{L_{i}}^{G}(V)) is a submodule of soc(Res_{L_{i}}^{G}(X)) for each i, so the result follows.

REMARKS: i) It need not be the case in part i) that X is projective as kG/C_{G}(X)-module. For example, if we consider the fusion system realised by M_{11} for p = 2, we are led to the amalgam G = GL(2, 3) ∗_{D_{8}} S_{4}. Take X_{1} to be the projective cover of the trivial module (over k, with p = 2) for L_{1} = GL(2, 3) and X_{2} to be the direct sum of two copies of the projective trivial module for L_{2} = S_{4}. As above, we may construct a module X for G via these two actions. Now G/C_{G}(X) is finite, but we claim that the image of P is not a Sylow 2subgroup of the factor group.

Suppose otherwise. We note that X may be afforded by a transitive permutation module, so that G/C_{G}(X) has a subgroup of index 16. In that case, for K a maximal free normal subgroup containing C_{G}(X), we see that H = G/K has a Hall 2′-subgroup and a semi-dihedral Sylow 2-subgroup Q of order 16. Hence every composition factor of H has a Hall 2′-subgroup, as does every section of H of order divisible by 16. We also know from earlier arguments that H has a unique minimal normal subgroup, which is perfect. Hence F^{∗}(H) is simple. If F^{∗}(H) does not have order divisible by 16, then it has a subgroup of index 8, so is isomorphic to a subgroup of A_{8}. Since it has a Hall 2′-subgroup, we have F^{∗}(H) ≅ PSL(2, 7). But Aut(PSL(2, 7)) has a dihedral Sylow 2-subgroup, so that F^{∗}(H)(= M, say) must contain Q. For the sake of brevity, we assume some knowledge of the structure of groups with a semi-dihedral Sylow 2-subgroup.

Setting C = C_{M}(t) for t a central involution of Q, we have C ≠ O^{2}(C), so that O^{2}(C)/⟨t⟩ has a Sylow 2-subgroup of order 4, and a subgroup of odd order of index 4. We must conclude that C/O(C) ≅ GL(2, 3), and that M ≅ M_{11} or PSL(3, 3). But neither of these candidates for M has a Hall 2′-subgroup, so we have arrived at a contradiction. Thus [G : C_{G}(X)] has order divisible by 32.

ii) If F is the fusion system of a finite group of Lie type in characteristic p and the L_{i}’s are the minimal parabolics containing a fixed Borel subgroup B, we may take (for each i) X_{i} to be the projective cover as kL_{i}-module of the non-trivial summand of the doubly transitive permutation module afforded by the action of L_{i} on the cosets of B. The module X constructed in part i) of Theorem 4 is not uniquely determined up to isomorphism by the isomorphism type of the X_{i}’s (uniqueness requires an explicit choice of action for each L_{i} with compatibility of the action of the P_{i}’s). However, one choice of the module X yields the Steinberg module for the original group of Lie type.

ACKNOWLEDGEMENT: We are indebted to R. Kessar and M. Linckelmann for carefully reading an earlier version of this paper and pointing out some textual corrections. We are also grateful to R. Stancu for pointing out an error in an earlier example used in Section 3.

REFERENCES

[1] Alperin,J.L., Sylow intersections and fusion, J. Algebra,6,(167),222-241.

[2] Alperin,J.L., Weights for Finite Groups, Proceedings of Symposia in Pure Mathematics 47, ( AMS, Providence,1987),369-379.

[3] Alperin,J.L.; Broué,M. Local Methods in Block Theory, Annals of Mathematics (2),110(1979), 143-157.

[4] Aschbacher,M. & Chermak,A., A group-theoretic approach to a family of 2-local finite groups constructed by Levi and Oliver, Preprint, August 2005.

[5] Broto,C.,Castellana,N.,Grodal,J.,Levi,R. & Oliver,R., Subgroup families controlling p-local finite groups, Proc. LMS (3),91,(2005),3,325-354.

[6] Broto,C.,Levi,R. & Oliver,R.,The homotopy theory of fusion systems, J.Amer. Math. Soc.,16,(2003),4,779-856.

[7] Broto,C.,Levi,R. & Oliver,R.,Homotopy equivalence of p-completed classifying spaces of finite groups, Invent. Math.,151,(2003),3,611-664.

[8] Dade, E.C., Counting Characters in Blocks,I, Invent. Math. 109, 1,(1992), 187-210.

[9] Goldschmidt,D., A conjugation family for finite groups, J. Algebra,16,(1970),138-142.

[10] Kessar,R;Linckelmann,M;Robinson,G.R, Local control in fusion systems of p-blocks of finite groups, J. Algebra 257,2,(2002),393-413.

[11] Külshammer, B.; Puig, L., Extensions of Nilpotent Blocks, Invent. Math,102,1,(1990), 17-71.

[12] Levi,R.;Oliver,R. Construction of 2-local finite groups of a type studied by Solomon and Benson, Geometry and Topology,6,(2002),917-990.

[13] Puig,L., Sur le contrôle de fusion des p-sous-groupes, C.R.Acad.Sci. Paris, Ser. A, 278,(1974),231-232.

[14] Robinson, G.R., Dade’s Projective Conjecture for p-Solvable Groups, J. Algebra, 229,(2000),234-248.

[15] Robinson,G.R., Weight conjectures for ordinary characters, J. Algebra,275,(2004),761-775.

[16] Robinson,G.R., Large character heights, Qd(p), and the Ordinary Weight Conjecture,to appear in J. Algebra.

[17] Serre,J-P., Trees, Springer-Verlag,Berlin 1980.